OpenStudy (celinegirl):

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold. (1-6 ARE DIFFERENT QUESTIONS ABOUT ABOVE ^^)

OpenStudy (celinegirl):

1. Change the equation to slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all your work. 2. Describe how you would graph this line using the slope-intercept method. Be sure to write using complete sentences. 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4. Graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. 5. Suppose Sal's total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different. 6. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (celinegirl):

here is the graph:

OpenStudy (arivillafana):

gah damn, i can see why no one has bothered yet. I'd hate to have to do this assignment.

OpenStudy (anonymous):

@arivillafana that's not really helpful, but I sort of see what you mean, @celinegirl , try to do the first thing, if you can, if you can't then I'll try to explain what in the world is going on

OpenStudy (celinegirl):

@sarenemun yeah I don't really understand... It's due next week so im trying to get it done so I don't have to worry about it later. I'll just try to get one done at a time

OpenStudy (jango_in_dtown):

Ok lets start..

OpenStudy (jango_in_dtown):

we have 2x+3y=1470. This gives y=(2/3)x+490 Hence slope= (2/3) and y-intercept is 490

OpenStudy (jango_in_dtown):

@celinegirl

OpenStudy (anonymous):

I can help I'm actually doing it right now

OpenStudy (anonymous):

Mine is due tomorrow

OpenStudy (anonymous):

What problem do you need?